OpenStudy (anonymous):

A client comes to you for investment advice on his $500,000 winnings from the lottery. He has been offered the following options by three different financial institutions and requests assistance to help understand which option would be the best for his investment. Option 1: 6% compounded interest quarterly for 5 years. Option 2: 8% compounded interest annually for 5 years.

OpenStudy (tkhunny):

Option 1: 6% compounded interest quarterly for 5 years. Option 2: 8% compounded interest annually for 5 years On the surface, we seem to be missing Option 3. A little deeper, one has to wonder why compounding would make any difference at all when comparing 6% to 8%! Finally, You could calculate the values: (1 + 0.06/4)^(4*5) = ?? (1 + 0.08)^5 = ??

OpenStudy (anonymous):

^ is for what??

OpenStudy (tkhunny):

^ means exponentiation. 3^2 = 9 4^3 = 64 25^(1/2) = \(\sqrt{25}\) = 5

OpenStudy (anonymous):

(1 + 0.06/4)^(4*5) = 1.34685500655

OpenStudy (tkhunny):

Good. And the other one?

OpenStudy (anonymous):

(1 + 0.08)^5 = 1.4693280768

OpenStudy (anonymous):

thankyou so much

OpenStudy (tkhunny):

Very good. Now, how about that third option?

OpenStudy (anonymous):

There wasn't a third option...

OpenStudy (anonymous):

Explain to the client the main differences between simple interest versus compound interest. That is the only other question

OpenStudy (tkhunny):

"...options by three different financial..."? Oh, well.